# Hardware-Aware Quantum Kernel Design Based on Graph Neural Networks

### v1

@@ Abstract
Quantum kernel methods have emerged as a promising direction in quantum machine learning (QML), offering a principled way to map classical data into high-dimensional quantum Hilbert spaces. While conceptually powerful, designing effective quantum kernels that adapt to both the target task and the constraints of near-term quantum hardware remains a nontrivial challenge. In this work, we propose HaQGNN, a hardware-aware quantum kernel design framework that integrates quantum device topology, noise characteristics, and graph neural networks (GNNs) to evaluate and select task-relevant quantum circuits. By predicting surrogate metrics related to fidelity and kernel performance, HaQGNN enables efficient circuit screening at scale. Feature selection is further incorporated to improve compatibility with limited-qubit systems and mitigate kernel degradation. Extensive experiments on three benchmark datasets, Credit Card, MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms existing quantum kernel baselines in terms of classification accuracy. Our work highlights the potential of learning-based and hardware-aware strategies for advancing practical quantum kernel design on NISQ devices.
@@ I Introduction
Quantum computing presents a compelling prospect for revolutionizing machine learning tasks by exploiting the fundamental properties of quantum mechanics, such as superposition, entanglement, and high-dimensional Hilbert spaces [1, 2]. Quantum machine learning (QML), as a important class of quantum algorithms tailored for the Noisy Intermediate-Scale Quantum (NISQ) era [3], has demonstrated practical applicability in diverse real-world tasks, including classification [4, 5, 6, 7, 8], generative modeling [9, 10, 11, 12], and learning physical systems [13, 14]. Among various QML approaches, quantum kernel methods have attracted increasing attention due to their solid theoretical foundation and interpretability [15, 16, 17]. By embedding classical data into quantum states through parameterized quantum circuits, quantum kernels define implicit feature maps capable of capturing complex nonlinear relationships beyond classical means.
@@
Despite recent demonstrations of quantum kernel learning on real quantum devices, realizing practical quantum advantage remains challenging, especially under the constraints of Noisy Intermediate-Scale Quantum (NISQ) devices [18, 19, 20]. In particular, unsuitable quantum circuit designs may lead to vanishing kernel similarities [21], reduced expressivity, and degraded generalization [22, 23], severely limiting the performance of quantum kernels. These drawbacks are exacerbated when the data dimensionality is high or when the circuit depth and qubit count exceed the noise tolerance of near-term devices. To mitigate these limitations, several works have investigated circuit-aware and data-aware quantum kernel optimization. Some methods focus on tuning variational parameters in fixed circuit layouts [24, 25, 26], while others explore structural search via evolutionary or Bayesian algorithms [27, 28, 29]. However, most existing approaches either ignore hardware constraints such as qubit connectivity and gate noise, or suffer from excessive computational costs in searching large quantum circuit spaces.
Figure 1: An overview of handware-aware quantum kernel design.
@@
To address these challenges, we propose a hardware-aware quantum kernel design framework based on Graph Neural Networks (HaQGNN), which efficiently integrates quantum hardware topology and device noise characteristics into the design of quantum feature maps, as illustrated in Fig. 1. The use of GNNs builds upon our prior work [30], where graph-based models were employed to predict specific properties of quantum circuits with high accuracy. Rather than formulating quantum kernel design as a joint optimization over circuit structure and parameters, our approach decouples the process by first generating hardware-aware candidate circuits and then employing graph neural networks to evaluate their potential performance. Leveraging the graph structure of quantum circuits, we employ GNNs to predict two surrogate metrics: the ”Probability of Successful Trials (PST)” [31], which reflects circuit fidelity under hardware noise, and kernel-target alignment (KTA), which correlates with downstream classification accuracy.
@@
The HaQGNN framework consists of four steps: (1) candidate circuit generation based on hardware topology and native gate sets; (2) PST prediction via GNNs-1 to filter out low-fidelity circuits; (3) KTA prediction via GNNs-2 to select high-performing kernels; and (4) parameter optimization of the top-$k$ circuits for final task-specific performance. Furthermore, we incorporate feature selection to reduce input dimensionality, mitigating vanishing similarity issues and enhancing compatibility with NISQ devices. To the best of our knowledge, this is the first work to propose a GNN-based, hardware-aware framework that evaluates and selects quantum kernel circuits by jointly considering hardware topology, noise characteristics, and circuit structure, followed by task-specific parameter optimization. Extensive simulations on IBM quantum devices demonstrate the superiority of HaQGNN over several state-of-the-art quantum kernel baselines across multiple classification tasks and device configurations.
@@
Our proposed HaQGNN framework offers several distinctive advantages that collectively enable efficient quantum kernel design for NISQ devices:
@@
    Hardware-awareness: HaQGNN explicitly incorporates real-device topology and noise information during both circuit generation and performance prediction, enhancing compatibility and circuit fidelity across different quantum hardware backends.
@@
    Two-stage GNN-based performance estimation: By employing two lightweight graph neural networks (GNNs-1 and GNNs-2), the framework accurately predicts fidelity (PST) and classification potential (KTA) of circuits with minimal simulation cost, achieving speedups of over 380× and 58000× compared to traditional metrics.
@@
    Noise-robust candidate selection: By filtering out low-fidelity circuits early in the pipeline, HaQGNN effectively mitigates the impact of gate errors and decoherence, leading to more reliable kernel performance on noisy devices.
@@
    Superior classification accuracy: Extensive experiments on multiple benchmark tasks (e.g., Credit Card, MNIST-5, FMNIST-4) and across various IBM quantum devices (e.g., Perth, Lagos, Torino) demonstrate that HaQGNN consistently outperforms several state-of-the-art quantum kernel baselines.
The remainder of this paper is organized as follows. Section II introduces background knowledge on quantum computing, kernel methods, and graph neural networks. Section III describes the proposed HaQGNN framework in detail. Section IV presents numerical results and comparative evaluations. Finally, Section V concludes the paper and discusses future research directions.
@@ II.3 Graph neural networks
Graph Neural Networks (GNNs) have become a powerful framework for learning over graph-structured data, where relationships among entities play a critical role in the underlying data distribution [32]. By iteratively aggregating and transforming information from neighboring nodes through message-passing mechanisms, GNNs are capable of capturing both local topology and global structural dependencies [33]. This capacity to model complex relational patterns has led to their widespread adoption across a variety of domains, including social network analysis [34], molecular property prediction [35, 36, 37], and recommender systems [38, 39]. Notably, variants such as Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and GraphSAGE have significantly advanced the state-of-the-art in graph representation learning tasks [40, 41, 42, 43, 44]. In the context of quantum computing, quantum circuits can be naturally represented as graphs, where qubits correspond to nodes and quantum gates or couplings define the edges. This structural analogy provides a compelling motivation for applying GNNs to quantum systems, enabling efficient learning and prediction tasks grounded in the intrinsic connectivity of quantum architectures.
@@ III Method
In this section, we present the technical details of our proposed hardware-aware quantum kernel design framework. An overview of the framework is illustrated in Fig. 1, which consists of four main steps:
Step 1: Based on the hardware topology and noise characteristics of the quantum device, a set of qubits is selected such that they exhibit favorable connectivity and minimal susceptibility to noise. A large number of candidate quantum circuits are then generated on the selected qubits, following the supported native gate set and qubit connectivity of the target device.
Step 2: A graph neural network (GNNs-1) is employed to predict the ”Probability of Successful Trials (PST)”, a metric correlated with circuit fidelity. This step enables the early rejection of low-fidelity candidate circuits, thereby filtering out circuits with potentially higher fidelity.
Step 3: A second graph neural network (GNNs-2) is designed to estimate the kernel-target alignment (KTA), a surrogate metric for classification accuracy, over the high-fidelity circuits identified in Step 2. The top-$k$ circuits with the highest predicted KTA values are selected for further refinement.
Step 4: A classical optimizer is employed to train the parameters of the top-$k$ high-KTA circuits with the high PST values from scratch, and the circuit with the best task-specific classification accuracy is selected as the final quantum kernel circuit.
We divide the hardware-aware quantum kernel design into three parts: candidate circuit generation (Step 1, Section III.1), metric prediction based on graph neural networks (Step 2-3, Section III.2), circuit training and selection (Step 4, Section III.3).
@@ III.1 Candidate circuit generation
The process of generating candidate quantum circuits is primarily divided into three stages: 1) A specific number of qubits with preserved topological connectivity are selected from the quantum device. This selection process takes into account the noise characteristics of individual qubits and their couplings, aiming to identify a subgraph of the device topology with the lowest average noise. Such the selection strategy improves hardware efficiency and enhances the fidelity of the generated circuits. 2) Candidate circuits are generated based on the selected qubit and the native gate set supported by the quantum device. During the circuit generation process, each inserted quantum gate must act on qubits that are physically connected on the hardware. This ensures that qubit mapping is implicitly achieved during circuit construction, thereby avoiding the insertion of costly SWAP gates during the compilation phase. As a result, the overall circuit depth is reduced, and noise accumulation is mitigated. 3) In the candidate quantum circuits, parameterized quantum gates are selected at different positions to enable flexible data embedding, which helps improve the performance of the quantum kernel.

### v2

# HaQGNN: Hardware-Aware Quantum Kernel Design Based on Graph Neural Networks

@@ Abstract
Designing effective quantum kernels is a central challenge in Quantum Machine Learning (QML), particularly under the limitations of Noisy Intermediate-Scale Quantum (NISQ) devices with a limited number of qubits, error-prone gate operations, and restricted qubit connectivity. To address this, we propose HaQGNN, a hardware-aware quantum kernel design method that integrates quantum device topology, noise characteristics, and Graph Neural Networks (GNNs) to evaluate and select task-relevant quantum circuits that define quantum kernels. First, each quantum circuit is represented as a directed acyclic graph that encodes hardware-specific features, including gate types, target qubits, and noise characteristics. Next, two GNNs are trained to predict surrogate metrics, Probability of Successful Trials (PST) and Kernel-Target Alignment (KTA), for fast and accurate fidelity and performance estimation. Additionally, feature selection is further incorporated to reduce input dimensionality and improve compatibility with limited-qubit devices. Finally, extensive experiments on three benchmark datasets, Credit Card (CC), MNIST-5, and FMNIST-4, demonstrate that HaQGNN outperforms existing baselines in terms of classification accuracy. Our results highlight the potential of learning-based and hardware-aware strategies for advancing practical quantum kernel design on near-term quantum hardware.
@@ I Introduction
Quantum computing presents a compelling prospect for revolutionizing machine learning tasks by exploiting the fundamental properties of quantum mechanics, such as superposition, entanglement, and high-dimensional Hilbert spaces [1, 2]. Quantum Machine Learning (QML), as an important class of quantum algorithms tailored for the Noisy Intermediate-Scale Quantum (NISQ) era [3], has demonstrated practical applicability in diverse real-world tasks, including classification [4, 5, 6, 7, 8], generative modeling [9, 10, 11, 12], and learning physical systems [13, 14]. Among various QML approaches, quantum kernel methods have attracted increasing attention due to their solid theoretical foundation and interpretability [15, 16, 17]. By embedding classical data into quantum states through parameterized quantum circuits, quantum kernels define implicit feature maps capable of capturing complex nonlinear relationships beyond classical means.
@@
Despite recent demonstrations of quantum kernel learning on real quantum devices, realizing practical quantum advantage remains challenging, especially under the constraints of NISQ devices [18, 19, 20]. In particular, unsuitable quantum circuit designs may lead to vanishing kernel similarities [21], reduced expressivity, and degraded generalization [22, 23], severely limiting the performance of quantum kernels. These drawbacks are exacerbated when the data dimensionality is high or when the circuit depth and qubit count exceed the noise tolerance of near-term devices. To mitigate these limitations, several works have investigated circuit-aware and data-aware quantum kernel optimization. Some methods focus on tuning variational parameters in fixed circuit layouts [24, 25, 26], while others explore structural search via evolutionary or Bayesian algorithms [27, 28, 29]. However, most existing approaches either ignore hardware constraints such as qubit connectivity and gate noise, or suffer from excessive computational costs in searching large quantum circuit spaces.
@@
To address these challenges, we propose HaQGNN, a hardware-aware quantum kernel design based on Graph Neural Networks, which efficiently integrates quantum hardware topology and device noise characteristics into the design of quantum feature maps, as illustrated in Fig. 1. The use of GNNs builds upon our prior work [30], where graph-based models were employed to predict specific properties of quantum circuits with high accuracy. Rather than formulating quantum kernel design as a joint optimization over circuit structure and parameters, our approach decouples the process by first generating hardware-aware candidate circuits and then employing graph neural networks to evaluate their potential performance. Leveraging the graph structure of quantum circuits, we employ GNNs to predict two surrogate metrics: the Probability of Successful Trials (PST) [31], which reflects circuit fidelity under hardware noise, and Kernel-Target Alignment (KTA), which correlates with downstream classification accuracy.
@@
The HaQGNN method consists of four steps: (1) candidate circuit generation based on hardware topology and native gate sets; (2) PST prediction via GNNs-1 to filter out low-fidelity circuits; (3) KTA prediction via GNNs-2 to select high-performing kernels; and (4) parameter optimization of the top-$k$ circuits for final kernel selection. Furthermore, we incorporate feature selection to reduce input dimensionality, mitigating vanishing similarity issues and enhancing compatibility with NISQ devices. To the best of our knowledge, this is the first work to propose a GNN-based, hardware-aware method that evaluates and selects quantum kernel circuits by jointly considering hardware topology, noise characteristics, and circuit structure, followed by task-specific parameter optimization. Extensive simulations on IBM quantum devices demonstrate the superiority of HaQGNN over several state-of-the-art quantum kernel baselines across multiple classification tasks and device configurations.
@@
Our proposed HaQGNN method offers several distinctive advantages that collectively enable efficient quantum kernel design for NISQ devices:
@@
    Hardware-awareness: HaQGNN explicitly incorporates real-device topology and noise information during both circuit generation and performance prediction, enhancing circuit performance across different quantum hardware backends.
@@
    Dual-GNN-based performance prediction: By employing two graph neural networks (GNNs-1 and GNNs-2), the HaQGNN accurately predicts fidelity (PST) and classification accuracy (KTA) of circuits with low simulation cost, achieving speedups of over 380× and 58000× compared to direct calculation.
@@
    Noise-aware candidate filtering: By filtering out low-fidelity circuits early in the pipeline, HaQGNN effectively mitigates the impact of gate errors and decoherence, leading to more reliable kernel performance on noisy devices.
@@
    Superior classification accuracy: Extensive experiments on multiple benchmark tasks (e.g., Credit Card (CC), MNIST-5, FMNIST-4) and across various IBM quantum devices (e.g., Perth, Lagos, Torino) demonstrate that HaQGNN consistently outperforms several state-of-the-art baselines.
The remainder of this paper is organized as follows. Section II introduces background knowledge on quantum computing, kernel methods, and graph neural networks. Section III describes the proposed HaQGNN method in detail. Section IV presents numerical results and comparative evaluations. Finally, Section V concludes the paper and discusses future research directions.
@@ II.3 Graph neural networks
Graph Neural Networks (GNNs) have become a powerful framework for learning over graph-structured data, where relationships among entities play a critical role in the underlying data distribution [32]. By iteratively aggregating and transforming information from neighboring nodes through message-passing mechanisms, GNNs are capable of capturing both local topology and global structural dependencies [33]. This capacity to model complex relational patterns has led to their widespread adoption across a variety of domains, including social network analysis [34], molecular property prediction [35, 36, 37, 38], and recommender systems [39, 40]. Notably, variants such as Graph Convolutional Networks (GCNs), Graph Attention Networks (GATs), and GraphSAGE have significantly advanced the state-of-the-art in graph representation learning tasks [41, 42, 43, 44, 45]. In the context of quantum computing, quantum circuits can be naturally represented as graphs, where qubits correspond to nodes and quantum gates or couplings define the edges. This structural analogy provides a compelling motivation for applying GNNs to quantum systems, enabling efficient learning and prediction tasks grounded in the intrinsic connectivity of quantum architectures.
@@ III Method
Figure 1: An overview of HaQGNN.
In this section, we present the technical details of our proposed HaQGNN. An overview of this method is illustrated in Fig. 1, which consists of four main steps:
Step 1 Candidate Circuit Generation: Based on the hardware topology and noise characteristics of the quantum device, a set of qubits is selected such that they exhibit favorable connectivity and low susceptibility to noise. A large number of candidate quantum circuits are then generated on the selected qubits, following the supported native gate set and qubit connectivity of the target device.
Step 2 PST Prediction: A graph neural network (GNNs-1) is employed to predict the PST, a metric correlated with circuit fidelity. This step enables the early rejection of low-fidelity candidate circuits, thereby filtering out circuits with potentially higher fidelity.
Step 3 KTA Prediction: A second graph neural network (GNNs-2) is designed to estimate the KTA, a surrogate metric for classification accuracy, over the high-fidelity circuits identified in Step 2. The top-$k$ circuits with the highest predicted KTA values are selected for further refinement.
Step 4 Final Training and Kernel Selection: A classical optimizer is employed to train the parameters of the top-$k$ high-KTA circuits with the high PST values from scratch, and the circuit with the best task-specific classification accuracy is selected as the final quantum kernel.
To facilitate presentation, we organize the four-step HaQGNN method into three parts: candidate circuit generation (Step 1, Section III.1), metrics prediction based on graph neural networks (Step 2-3, Section III.2), final training and kernel selection (Step 4, Section III.3).
@@ III.1 Candidate circuit generation
The process of generating candidate quantum circuits is primarily divided into three stages: 1) A specific number of qubits with preserved topological connectivity are selected from the quantum device. This selection process takes into account the noise characteristics of individual qubits and their couplings, aiming to identify a subgraph of the device topology with the lowest average noise. Such a selection strategy improves hardware efficiency and enhances the fidelity of the generated circuits. 2) Candidate circuits are generated based on the selected qubit and the native gate set supported by the quantum device. During the circuit generation process, each inserted quantum gate must act on qubits that are physically connected on the hardware. This ensures that qubit mapping is implicitly achieved during circuit construction, thereby avoiding the insertion of costly SWAP gates during the compilation phase. As a result, the overall circuit depth is reduced, and noise accumulation is mitigated. 3) In the candidate quantum circuits, parameterized quantum gates are selected at different positions to enable flexible data embedding, which helps improve the performance of the quantum kernel.